You have accumulated $4,100 and are looking for the best rate of return that can be earned over the next year. A bank savings account will pay 6%. A one-year bank certificate of deposit will pay 8%, but the minimum investment is $7,100.

a. Calculate the amount of return you would earn if the $4,100 were invested for one year at 6%.

b. Calculate the net amount of return you would earn if $3,000 were borrowed at a cost of 16%, and then $7,100 were invested for one year at 8%. (Round intermediate calculations.)

c. Calculate the net rate of return on your investment of $4,100 if you accept the strategy of part b. (Round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

a. Amount of return = $4,100 * 6% = $246

b. Amount of return = $7,100 * 8% = $568

Borrowing cost = $3,000 * 16% = $480

Net amount of return = Amount of return - Borrowing cost = $568 - $480 = $88

c. Net rate of return = [ $88 / $4,100 ] * 100 = 2.15%